A car travels 89 miles east in 22 minutes. What is the velocity in miles per hour? You do not need to include the units in the answer but you must include the direction.
Answer by jim_thompson5910(35256)

60 min = 1 hr

1 min = 1/60 hr

22 min = 22/60 hr

22 min = 11/30 hr

Velocity = 89/(11/30) = 2670/11 = 242.73 mph east